The perimeter of a rectangle is 76 feet. The lenght is three feet less than twice the width. Find the length and with.
Answer by JulietG(1812)

Remember that a perimeter is the distance around -- think of a perimeter fence. How much do you need to fence your property (the rectangle)? The formula is 2W + 2P = P
.
We're given the following:
2W + 2L = 76
L = 2W - 3
.
Substitute that value for L to get:
2W + 2(2W-3) = 76
Distribute:
2W + 4W - 6 = 76
Add 6 to each side:
2W + 4W = 82
Add the Ws:
6W = 82
Divide each side by 6:
W = 13 2/3
Since L = 2W - 3, L = 27 1/3 - 3 = 24 1/3
.
That's an odd choice. Let's plug it in to make certain.
2W + 2L = 76
2(13 2/3) + 2(24 1/3) = 76
27 1/3 + 48 2/3 = 76
That indeed is correct. It's still an odd number choice...